1.     a. Initially, the demand and supply for beans are


where p = price in cents per pound and Q = pounds per day. The government has a stockpile of beans (which is not included in the initial supply equation). It wants to cut the price of beans to 8 cents per pound. How much should it sell from its stockpile?

 

b. Explain each of the following cases.

(a)   If the elasticity is greater than 1, is demand elastic or inelastic? If the elasticity equals 0, is demand perfectly elastic or perfectly inelastic?

(b)  If demand is elastic, how will an increase in price change total revenue?



1
Expert's answer
2021-10-13T10:17:38-0400

1)

"Qd=Qs\\100000-50000p= -50000+10000p\\15000p=150000\\p=10\\Q=100000-50000(8)\\p=50000units"

If government wants to cut price to 8 then

"Qd=100000-5000(8)\\=60000

\\

Qs=-50000-10000(8)\\=30000\\Qd>Qs"

The government needs to sell 60000-30000=30000 from its stock pile.

b)

If elasticity is greater than 1 demand is elastic. If elasticity is equal to zero , demand is perfectly inelastic.

b)

Increase in price Will reduce total revenue.